Use the following financial statement information available for Jones Corporation

      2012                           2011     

Net sales                    $784,000                     $697,000

Cost of goods sold       406,000                       377,000

Net income                   112,000                         80,000

Tax expense                  48,000                         29,000

Interest expense            14,000                         14,000

(a) The net profit margin ratio for Jones Corporation for 2012 is:

(b) The net profit margin ratio for Jones Corporation for 2011 is:

(c) The gross profit margin ratio for Jones Corporation for 2012 is:
Answer #1
Calculations Particulars 2012 2011
A Net Sales 784000 697000
B Cost of goods sold 406000 377000
C = A - B Gross Profit 378000 320000
D Net Income 112000 80000
E = D / A x 100 Net profit margin 14.29% 11.48%
F = C / A x 100 Gross Profit margin 48.21% 45.91%

(a) The net profit margin ratio for Jones Corporation for 2012 is: 14.29%

(b) The net profit margin ratio for Jones Corporation for 2011 is: 11.48%

(c) The gross profit margin ratio for Jones Corporation for 2012 is: 48.21%